# CrowdSurfer: Sampling Optimization Augmented with Vector-Quantized Variational AutoEncoder for Dense Crowd Navigation

### v1

@@ IV-E Adapting to Changing Map
In this subsection, we test the adaptivity of our approach and DRL-VO to the changes in the static environment. We take the Square World environment with 25 dynamic pedestrians and add additional static obstacles. These obstacles were not seen during the mapping and thus, the prior-compute plan becomes infeasible on many runs. Table II summarizes the key results for 9 trials. As can be seen, the DRL-VO success rate dropped from 0.75 (Table I) to 0.44. In contrast, our approach showed a very marginal drop to 0.89 from 0.90 (Table I).
@@ IV-F Reliance on Global Plan
Table I (second half) compares the performance of DRL-VO and our approach in the absence of guidance from the global plan. More specifically, the global plan is simply a straight line connecting the start and the goal. As can be seen, although the success rate reduces for both approaches, the performance falloff is steeper for DRL-VO dropping approximately 28% on average. In contrast, our approach is much more robust dropping approximately 15% on average.
@@ IV-G Ablation: Comparison with PRIEST [1]
Our pipeline builds on PRIEST by augmenting it with a learned and environment-conditioned initialization through the VQ-VAE+PixelCNN model. Thus, this sub-section analyzes the impact that our learned model provides over the regular PRIEST pipeline. We consider the Cumberland environment with 35 dynamic agents as PREIST performed best in this setting. The results are summarized in Table III. We showcase the statistics for different computation budgets of PRIEST (2, 10, 12 iterations). As can be seen, the environment conditioned prior provided by VQ-VAE+PixelCNN dramatically improves the success rate along with average travel time, path length, and speed.
@@ IV-H Real World Configuration
The testing was done in the lab at RRC, IIIT Hyderabad, and various static and dynamic obstacles were added during navigation, as can be seen in Figure LABEL:fig:fig_teaser.
@@ V Conclusion
This paper proposed a novel framework that couples generative prior models learned on expert data with inference time optimization to show results that are significantly better than prior art in highly populated pedestrian worlds containing as many as 55 moving pedestrians. While the generative priors are efficient in exploring the homotopies and are inherently multimodal they are not goal-conditioned and are not guaranteed to avoid collisions. The method of inference time batch projection ensures both these conditions are met with a very high degree of accuracy. Further, the efficacy of the proposed framework in the absence of a global plan or in changed environments where the planner is rendered ineffective is a cornerstone of this effort. In the future, we expect to condition the generative prior on both image and range data even as we focus on improving real-time performance with higher fidelity.
@@ Acknowledgements
The author, Laksh Nanwani, thanks IHub-Data, IIIT Hyderabad, for extending their research fellowship. We also acknowledge IHub-Data for supporting this work.

### v2

@@ IV-E Adapting to Changing Map
In this subsection, we test the adaptivity of our approach and DRL-VO to changes in the static environment. We take the Square World environment with 25 dynamic pedestrians and add additional static obstacles. These obstacles were not seen during the mapping, and thus, the prior-compute plan became infeasible on many runs. Table II summarizes the key results for 9 trials. As can be seen, the success rate of DRL-VO dropped from 0.80 (Table I) to 0.44. In contrast, our approach showed a marginal drop to 0.89 from 0.90 (Table I).
@@ IV-F Reliance on Global Plan
Table I (second half) compares the performance of DRL-VO and our approach without guidance from the global plan. More specifically, the global plan is just a straight line connecting the start and the goal. As can be seen, although the success rate reduces for both approaches, the performance falloff is steeper for DRL-VO, dropping approximately 28% on average. In contrast, our approach is much more robust, dropping approximately 15% on average.
@@ IV-G Ablation: Comparison with PRIEST [1]
Our pipeline builds on PRIEST by augmenting it with a learned and environment-conditioned initialization through the VQ-VAE+PixelCNN model. Thus, this sub-section analyzes our learned model’s impact over the regular PRIEST pipeline. We consider the Cumberland environment with 35 dynamic agents as PRIEST performed best in this setting. The results are summarized in Table III. We showcase the statistics for different computation budgets of PRIEST (2, 10, 12 iterations). As can be seen, the environment-conditioned priors provided by VQ-VAE+PixelCNN dramatically improve the success rate along with average travel time and speed.
@@ IV-H Real World Configuration
The testing was done at RRC, IIIT Hyderabad, as shown in Figure 1. We were able to achieve reasonable success rates both indoors and outdoors, albeit lower than the simulation testing. One of the many reasons could be inaccuracies in the human-detection pipeline, which can be improved in the future using more sophisticated methods. While 2 CEM iterations delivered reasonable success rates, increasing the number of iterations to 3 achieved a better success rate in the real world.
@@ V Conclusion
This paper proposed a novel framework that couples generative prior models learned on expert data with inference-time optimization to show results that are significantly better than prior art in highly populated pedestrian worlds containing as many as 55 moving pedestrians. While the generative priors are efficient in exploring the homotopies and are inherently multimodal, they are not goal-conditioned and are not guaranteed to avoid collisions. The method of inference-time batch projection ensures both these conditions are met with a high degree of accuracy. Further, the efficacy of the proposed framework in the absence of a global plan or in changed environments where the planner is rendered ineffective is a cornerstone of this effort. In the future, we expect to condition the generative prior on both image and range data even as we focus on improving real-time performance with higher fidelity.
@@ Acknowledgements
The author, Laksh Nanwani, thanks IHub-Data, IIIT Hyderabad, for extending their research fellowship. We also acknowledge IHub-Data for supporting this work via project M2-029. This work was also funded in part by the European Social Fund and Estonian Research Council via project TEM-TA101 and Grant PSG753 from the Estonian Research Council.